Suppose a firm operating in a perfectly competitive industry has costs in the short run given by:

SRTC = 8 + 1/2Q^2 and therefore MC q.



If the minimum point of the short-run ATC curve for all firms(existing and potential)is also the minimum point of the long-run average cost curve (LRAC), calculate the long-run equilibrium price, market quantity, and firm quantity. What is the long-run equilibrium number of firms in the industry?


1
Expert's answer
2021-11-14T17:58:27-0500



Short-run total cost = "8+1\/2Q^2"

MC = Q

Let's say Demand function = 1000 - 100P for instance, then

1) The short-run supply curve for the firm is the portion of the marginal cost curve that is above the average variable cost curve.


Short-run supply curve of a competitive firm is given by

P = MC

P = Q ...( Short-run supply equation)


2) Let us assume there are 100 identical firms ,

the supply curve of the market is given by , Qs = nQ= 100*P = 100P

3) At equilibrium Qs= Qd

100P = 1000 -100P

200P = 1000

P = 5

put this value in either demand or supply function

Q= 1000 - 5*100

= 1000 - 500

= 500

Equilibrium quantity of the market is 500

Firm Quantity = Market quantity / 100

"=\frac{ 500}{100}"

= 5


4) Profits are ZERO in the competitive market in the long-run ..

With the available prices and quantity the profit is not zero in this case.

so it is not the long-run equilibrium.